OpenStudy (96mertcank96):

express the interval in set builder notation and graph the interval on a number line (-1,1]

OpenStudy (irishboy123):

so you are telling us that that \(x\gt -1 \) and \( \ x\le 1\) or \(-1 \lt x \le 1\) so you might wish to start out with: \(\large \{ x : \ x\gt -1 \) and \( \ x\le 1 \ \}\) or \(\large \{ x \in \mathcal{Z} : \ x\gt -1 \) and \( \ x\le 1 \ \}\) or \(\large \{ x \in \mathcal{R} : \ x\gt -1 \) and \( \ x\le 1 \ \}\) or \(\large \{ x \in \mathcal{C} : \ x\gt -1 \) and \( \ x\le 1 \ \}\)

OpenStudy (irishboy123):

but i reckon that the number line is the interesting bit, in the short term!